Equation of tangent of the parabola y=x^2 in terms of slope

Grade:

1 Answers

grijesh
5 years ago
Here the equation of tangent is y=mx +m 2/4 where the line of tangent is y =mx+c 
For proof put the value of y from line of tangent to parabola i.e x 2=mx+c then put discrimant to 0.then u get c=m 2/4.